A manager receives a monthly salary of $2900.00. Use the tax rates of 6.2% for social security and 1.45% for medicare. Find the social securty tax and the medicare tax to be deducted from one month's paycheck
Answer by checkley77(12844)

(.062+.0145)*2,900=.0765*2,900=$221.85 ans. for the amount deducted.
.062*2,900=$179.80 for SS.
.0145*2,900=$42.05 for Medicare.